How much time will it take a car travelling at 88 km/hr (55 mi/hr) to travel 500km?
Found 2 solutions by solver91311, fractalier:
Answer by solver91311(24713)



Distance equals rate times time, so time equals distance divided by rate.

John

My calculator said it, I believe it, that settles it

Answer by fractalier(6550)

Since distance equals rate times time, or d = rt, we can find time t by
t = d/r = 500/88 = 5.68 hours